# FakeNewsGPT4: Advancing Multimodal Fake News Detection through Knowledge-Augmented LVLMs

### v1

@@ 3.1.2 Dual-branch Fine-grained Verification
In addition to extracting semantic correlations between visual and textual features, mining artifact traces is also crucial for distinguishing fake news. Unlike [CLS] token, visual patch tokens with position encoding Vaswani et al. (2017) contain richer local spatial information. Given visual patch tokens $u_{v}^{\rm pat}$, we propose a dual-branch fine-grained verification module to encode them into artifact embeddings, guided by grounding annotations.
@@
As depicted in Fig. 3-(b), the upper branch which comprises a text encoder and a pixel decoder, is designed to utilize language-driven contrastive learning for pixel-wise verification. Dense features with good language alignment can provide complementary benefits for fine-grained segmentation Jeong et al. (2023); Gu et al. (2023). Specifically, two class prompts are initially encoded by the text encoder to obtain corresponding features $F_{p}^{i}\left(i=1,2\right)$, representing “natural” and “unnatura” states. Detailed contrastive class prompts are presented in the Appendix A. To restore local spatial details, the pixel decoder with consecutive deconvolution layers converts low-resolution features $u_{v}^{\rm pat}$ into high-resolution features $F_{h}$. Since patch-level features are not aligned with the textual space, we project both textual features $F_{p}$ and visual features $F_{h}$ into a shared representation space. Subsequently, the projected features $\tilde{F}_{p}$ and $\tilde{F}_{h}$ are employed to compute similarity scores followed by upsampling:
$w^{i}=\text{Upsample}\left(\tilde{F}_{p}^{T}\tilde{F}_{h}\right),\,\,\left(i=1,2\right).$ (5)
@@
By applying the scores spatially, we can establish the manipulated segmentation map to achieve per-pixel prediction:
$M_{s}=\sum_{i=1}^{2}{\log\left(\frac{e^{w^{i}}}{\sum{e^{w^{i}}}}\right)}.$ (6)
@@
To enrich the representation of artifact traces, the lower branch employs multi-head attention for patch-level verification. We utilize a learnable token $q_{\mathrm{tok}}$ serving as a query, while visual features $u_{v}^{\rm pat}$ act as both the key and value. Through the multi-head attention mechanism, local information related to artifacts is aggregated within $u_{\rm agg}$ under the supervision of bounding box grounding:
$u_{{\rm agg}}=\mathrm{Attention}\left(q_{{\rm tok}},\tilde{u}_{v}^{\rm pat},\tilde{u}_{v}^{\rm pat}\right).$ (7)
@@
To leverage the artifact knowledge contained in the two-branch features (i.e., the manipulated segmentation map $M_{s}$ and the aggregated token $u_{{\rm agg}}$), we separately devise a pixel encoder with multiple convolution layers and a simple linear layer. In this manner, both $M_{s}$ and $u_{{\rm agg}}$ are converted into artifact embeddings aligned with the final vision-language model.
3.1.3 Forgery-aware Vision Language Model
@@
Due to the lack of instruction-follow data in the MFND task, we carefully design an instruction template following the conversational format of the Vicuna Chiang et al. (2023), as shown below:
###Human: <Img><ImageFeature></Img>
<ForgeryFeature>[Human Prompt] ###Assistant:
In this prompt, <ImageFeature> represents the visual tokens produced by the image encoder and <ForgeryFeature> is the combination of semantic and artifact embeddings. The human prompt adopts the multiple choice question answering format, as shown in the dialog box of Fig. 3-(c).
To unleash the potential knowledge of LVLMs in solving MFND tasks, we devise two prompt strategies to serve as more informative inputs in Fig. 3-(c). On the one hand, we utilize candidate answer heuristics Robinson and Wingate (2022); Shao et al. (2023b) to present both the question and answer options to LVLMs and make them predict the symbol (e.g., “A”) associated with the selected answer. This approach enables the language models to explicitly compare different candidate answers showcasing more accurate responses. On the other hand, we implement soft prompt tuning to introduce learnable continuous vectors while freezing the language models. These vectors combined with semantic embeddings facilitate the extraction of additional semantic information. Meanwhile, this approach reduces the burden of LVLMs to learn forgery-aware alignment, thereby mitigating the catastrophic forgetting problem.
3.1.4 Loss Function
Two groups of loss functions are employed in our training procedure: cross-entropy loss, and two-branch verification losses.
Cross-entropy Loss.
@@ 3.1.4 Loss Function
$\mathcal{L}_{\mathrm{ce}}=-\sum_{i=1}^{n}{y_{i}{\mathrm{log}}(p_{i})},$ (8)
@@
Dual-branch Verification Loss.
Two-branch verification losses are designed for the precise encoding of artifact traces guided separately by pixel-wise and patch-level verification. Pixel-wise verification introduces focal loss Lin et al. (2017) and dice loss Milletari et al. (2016) to facilitate supervision for the manipulated segmentation map $M_{s}$: $\mathcal{L}_{\mathrm{pixel}}=\mathcal{L}_{\mathrm{focal}}+\mathcal{L}_{\mathrm{dice}}$. Patch-level verification involves regressing the final bounding box by utilizing the aggregated token $u_{{\rm agg}}$ and computing the regression losses with the ground-truth box by introducing L1 loss and GIoU loss Rezatofighi et al. (2019): $\mathcal{L}_{\text{patch}}=\mathcal{L}_{\mathrm{L}_{1}}+\mathcal{L}_{\text{giou}}$. More details about verification losses are provided in Appendix B.
@@
$\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{pixel}}+\mathcal{L}_{\mathrm{patch}}.$ (9)
4 Experiments
@@
[…]
@@
FakeNewsGPT4 78.56 28.52 73.47 81.97 25.31 76.29 87.07 21.19 79.06 87.94 19.81 80.16
[…]
4.1 Experimental Setup
Dataset.
[…]
Evaluation Metrics.
We treat the multimodal fake news detection problem as a binary classification task. Following previous works Shao et al. (2023a); Nan et al. (2022), we apply the Area Under the Receiver Operating Characteristic Curve (AUC), Equal Error Rate (EER), the Accuracy Score (ACC) as our evaluation metrics.
Baselines.
The proposed FakeNewsGPT4 is compared with the following strong baseline models: 1) PandaGPT Su et al. (2023): The vanilla PandaGPT model effectively aligns visual features with the text space of LLMs, enabling it to perform complex multimodal tasks in a zero-shot manner. 2) PandaGPT+SPT: This model integrates PandaGPT with soft prompt tuning Lester et al. (2021) by using the predefined question prompt along with learnable continuous vectors to fine-tune the LVLM during the instruction tuning phase. 3) HAMMER Shao et al. (2023a): HAMMER employs two unimodal encoders to encode image and text embeddings with contrastive learning for alignment, and then summarize multimodal information through the multimodal aggregator.
@@ 4.1 Experimental Setup
FakeNewsGPT4 75.17 33.35 71.42 75.15 32.85 70.52 81.06 26.21 75.26 85.88 22.41 78.87
@@
Implementation Details.
We utilize the visual encoder and text encoder sourced from ImageBind-Huge Girdhar et al. (2023) as the backbone for our image and text feature extractors. Moreover, we employ Vicuna-7B Chiang et al. (2023) as the inferential LLM, connected through a linear layer. The model is initialized from the instruction-tuned checkpoint provided by PandaGPT. All training images are resized to 224 × 224 and subjected to random horizontal flipping, as well as random perturbation techniques such as JPEG compression and Gaussian Blur following Shao et al. (2023a). The multi-level cross-modal reasoning module extracts intermediate patch features from the 8th, 16th, 24th, and 32nd layers of the image encoder. We set the learning rate as 1.5e-5 with a batch size of 16 and a maximum of 10 epochs when trained on the BBC subset. Linear warm-up and the one-cycle cosine learning schedule are adopted. All experiments are conducted on four NVIDIA GeForce 3090 GPUs with PyTorch. More details about hyperparameter settings are provided in Appendix D.
4.2 Performance Comparison
We evaluate the cross-domain performance of our FakeNewsGPT4 with baselines in both single-domain and multiple-domain settings respectively.
Single-domain Setting.
Table 2 presents the performance of our method and other baseline models in the challenging scenario where a single domain is available. We randomly select one subset as the source domain for training and the remaining subsets as the target domains for testing. From the results, we make the following observations: 1) The performance of the existing MFND method drops significantly when tested on the unknown subsets, which verifies the existence of domain shift caused by the deviation in propagation contents. 2) FakeNewsGPT4 yields substantial improvement on recent LVLMs, PandaGPT, and PandaGPT with soft prompt tuning, in both intra-domain and cross-domain testing. Such huge improvement demonstrates the effectiveness of forgery-specific knowledge augmentation in our framework. 3) Compared with the state-of-the-art method, HAMMER, our approach shows more remarkable improvement in cross-domain testing. For instance, for models trained on the BBC subset, FakeNewsGPT4 achieves a 7.7% increase in AUC when testing on the Washington Post subset. This may be credited to the effective utilization of inherent world knowledge from LVLMs in mitigating distribution discrepancies. The combination of forgery-specific knowledge and world knowledge facilitates profound manipulation reasoning in FakeNewsGPT4.
Multiple-domain Setting.
The inclusion of domestic news such as the BBC and The Guardian from British, as well as USA TODAY and The Washington Post from America, increases dataset diversity in practical scenarios. We select two subsets from identical countries for training and the remaining two for testing. The results are summarized in Table 3. Our FakeNewsGPT4 exhibits significant superiority over both PandaGPT using soft prompt tuning and HAMMER by a large margin. This reveals the effectiveness of our framework in instance-wise domain generalization guided by world knowledge derived from LVLMs, even when jointly learning multiple source domains.
@@ 4.2 Performance Comparison
ML&CM Reasoning DB&FG Verification Guardian USA Wash. Avg.
@@
Table 4: Ablation study of component modules. We evaluate the AUC (in %), EER (in %), and ACC (in %) of variant models on the remaining three subsets when trained on the BBC subset. ML: the extraction of multi-level patch features in the cross-modal (CM) reasoning module. DB: the extraction of dual-branch artifact features in the fine-grained (FG) verification module. Avg. denotes the mean value on the three testing subsets.
@@
FakeNewsGPT4 84.95 73.10 74.81 77.62
Table 5: Ablation study of prompt strategies. AUC (in %) of variant models is reported on the remaining three subsets when trained on the BBC subset. SPT denotes soft prompt tuning, whereas CAH refers to candidate answer heuristics.
4.3 Ablation Study
We perform several ablation experiments to explore the necessity of the proposed component modules, prompt strategies, and LVLMs knowledge, respectively. In the following experiments, all ablation results are evaluated on the remaining three subsets when trained on the BBC subset.
The Effect of the Component Modules.
In Table 4, we conduct a comprehensive ablation study on the proposed component modules to verify their effectiveness. The first row of Table 4 shows our baseline model that only performs soft prompt tuning, achieving an average AUC of 52.5%. Based on this baseline, we further introduce two separate modules: a multi-level cross-modal (ML&CM) reasoning module and a dual-branch fine-grained verification (DB&FG) module, with 21.55% and 20.16% improvement in average AUC, respectively. This implies that forgery-specific knowledge augmentation is indispensable for our framework. Comparatively, our model with complete two modules obtains the best performance increasing by 25.12%, indicating the effectiveness and complementarity of these two modules. Moreover, we test the performance of FakeNewsGPT4 removing the multi-level features (w/o ML), and FakeNewsGPT4 removing any one of the dual-branch features (w/o DB). These variant models lack rich features to represent forgery-specific knowledge leading to a great decrease in cross-domain performance.
The Effect of the Prompt Strategies.

### v2

@@ 3.3. Dual-branch Visual-artifact Localization
In addition to extracting semantic correlations between visual and textual features, mining artifact traces is also crucial for distinguishing fake news. Unlike [CLS] token, visual patch tokens with position encoding (Vaswani et al., 2017) contain richer local spatial information. Given visual patch tokens $u_{v}^{\rm pat}$, we propose a visual-artifact localization module to encode them into artifact embeddings, guided by grounding annotations.
@@
As depicted in Fig. 2-(b), the upper branch which comprises a text encoder and a pixel decoder, is designed to utilize language-driven contrastive learning for pixel-wise localization. Dense features with good language alignment can provide complementary benefits for fine-grained segmentation (Jeong et al., 2023; Gu et al., 2023). Specifically, two class prompts are initially encoded by the text encoder to obtain corresponding features $F_{p}^{i}\left(i=1,2\right)\in\mathbb{R}^{2\times C_{text}}$, representing “natural” and “unnatural” states. Detailed contrastive class prompts are presented in the Supplementary Materials. To restore local spatial details, the pixel decoder with consecutive deconvolution layers converts low-resolution features $u_{v}^{\rm pat}\in\mathbb{R}^{hw\times C_{img}}$ into high-resolution features $F_{h}\in\mathbb{R}^{H\times W\times C_{img}}$. Since patch-level features are not aligned with the textual space, we project both textual features $F_{p}$ and visual features $F_{h}$ into a shared representation space. Subsequently, the projected features $\tilde{F}_{p}\in\mathbb{R}^{2\times C}$ and $\tilde{F}_{h}\in\mathbb{R}^{H\times W\times C}$ are employed to compute similarity scores $w\in\mathbb{R}^{H\times W\times 2}$:
(5) $w^{i}=\tilde{F}_{h}\cdot\tilde{F}_{p}^{T},\,\,\left(i=1,2\right).$
@@
By applying the scores $w$ spatially, we can establish the manipulated segmentation map $M_{s}\in\mathbb{R}^{H\times W\times 2}$ to achieve per-pixel prediction:
(6) $M_{s}^{j}=\text{softmax}\left(w\right)=\log\left(\frac{e^{w^{j}}}{\sum_{i=1}^{2}{e^{w^{i}}}}\right),\,\,\left(j=1,2\right).$
@@
To enrich the representation of artifact traces with multiple levels of details, the lower branch employs multi-head attention for patch-level localization. We utilize a learnable token $q_{\mathrm{tok}}\in\mathbb{R}^{1\times C_{img}}$ serving as a query, while visual features $u_{v}^{\rm pat}$ act as both the key and value. Through the multi-head attention mechanism, local information related to artifacts is aggregated within $u_{\rm agg}\in\mathbb{R}^{1\times C_{img}}$ under the supervision of bounding box grounding:
(7) $u_{{\rm agg}}=\mathrm{Attention}\left(q_{{\rm tok}},\tilde{u}_{v}^{\rm pat},\tilde{u}_{v}^{\rm pat}\right).$
@@
To leverage the artifact knowledge contained in the two-branch features (i.e., the manipulated segmentation map $M_{s}$ and the aggregated token $u_{{\rm agg}}$), we separately devise multiple convolution layers and a simple linear layer as projectors. In this manner, both $M_{s}$ and $u_{{\rm agg}}$ are converted into continuous artifact embeddings aligned with the final vision-language model.
3.4. Forgery-aware Vision Language Model
@@
Due to the lack of instruction-follow data in the MFND task, we carefully design an instruction template following the conversational format of the Vicuna model (Chiang et al., 2023), as shown below:
###Human: <Img><ImageFeature></Img><ForgeryFeature>[Human Prompt] ###Assistant:
In this prompt, ¡ImageFeature¿ represents the visual tokens produced by the image encoder and ¡ForgeryFeature¿ is the combination of semantic and artifact embeddings. The human prompt adopts the multiple choice question answering format, as shown in the dialog box of Fig. 2-(c).
To unleash the potential knowledge of LVLMs in solving MFND tasks, we devise two prompt strategies to serve as more informative inputs in Fig. 2-(c). On the one hand, we utilize candidate answer heuristics (Robinson and Wingate, 2022; Shao et al., 2023b) to present both the question and answer options to LVLMs and make them predict the symbol (e.g., “A”) associated with the selected answer. This approach enables the language models to explicitly compare different candidate answers showcasing more accurate responses. On the other hand, we implement soft prompt tuning to introduce learnable continuous vectors while freezing the language models. These vectors combined with semantic embeddings facilitate the extraction of additional semantic information. Meanwhile, this approach reduces the burden of LVLMs to learn forgery-aware alignment, thereby mitigating the catastrophic forgetting problem.
3.5. Loss Function
Two groups of loss functions are employed in our training procedure: cross-entropy loss, and two-branch localization losses.
3.5.1. Cross-entropy Loss.
@@ 3.5. Loss Function
(8) $\mathcal{L}_{\mathrm{ce}}=-\sum_{i=1}^{n}{y_{i}{\mathrm{log}}(p_{i})},$
@@
3.5.2. Dual-branch Localization Loss.
Two-branch localization losses are designed for the precise encoding of artifact traces guided by pixel-wise and patch-level localization, respectively. Pixel-wise localization introduces focal loss (Lin et al., 2017) and dice loss (Milletari et al., 2016) to enable supervision for the manipulated segmentation map $M_{s}$: $\mathcal{L}_{\mathrm{pixel}}=\mathcal{L}_{\mathrm{focal}}+\mathcal{L}_{\mathrm{dice}}$. Patch-level localization involves regressing the final bounding box with the aggregated token $u_{{\rm agg}}$ and computing the regression losses with the ground-truth box by introducing L1 loss and GIoU loss (Rezatofighi et al., 2019): $\mathcal{L}_{\text{patch}}=\mathcal{L}_{\mathrm{L}_{1}}+\mathcal{L}_{\text{giou}}$. More details about localization losses are provided in the Supplementary Materials.
@@
(9) $\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\mathrm{pixel}}+\mathcal{L}_{\mathrm{patch}}.$
4. Experiments
@@
Table 1. The statistics of the four subsets within the DGM4 dataset categorized by the news sources.
@@
Table 2. Single-domain performance (%) comparison of baseline models on DGM4 dataset. Specifically, we use one subset for training and the remaining subsets for testing. SPT denotes the utilization of soft prompt tuning. ( ) indicates the intra-domain performance. The better results in each group are in boldface.
@@
FKA-Owl 78.56 28.52 73.47 81.97 25.31 76.29 87.07 21.19 79.06 87.94 19.81 80.16
4.1. Experimental Setup
4.1.1. Dataset.
We evaluate the proposed method on DGM4 dataset [^1] https://github.com/rshaojimmy/MultiModal-DeepFake (Shao et al., 2023a) and NewsCLIPpings dataset [^2] https://github.com/g-luo/news˙clippings (Luo et al., 2021a).
DGM4. DGM4 dataset is the recently released large-scale multimodal manipulation dataset which comprises 230K image-text paired samples with over 77K pristine pairs and 152K manipulated pairs. In the DGM4 dataset, image manipulation involves face swapping and facial emotion editing while text manipulation includes sentence replacement and textual sentiment editing. The construction of the DGM4 dataset is based on the VisualNews dataset (Liu et al., 2021b), which is collected from multiple news agencies. Different agencies cover diverse regional perspectives, thematic focus, and language styles (see Supplementary Materials for the analysis of word clouds), resulting in substantial distribution discrepancies. To simulate the open-world domain-shift scenarios, we partition the DGM4 dataset into four subsets based on news sources: BBC, The Guardian, USA TODAY (USA), and The Washington Post (Wash.). The statistics of four subsets are listed in Table 1.
NewsCLIPpings. The NewsCLIPpings dataset is the largest real-world multimodal fake news detection benchmark. Forgery samples within this dataset are automatically created by retrieving out-of-context images that appear convincing for a given caption yet display inconsistencies in entities or semantic context. We follow the official dataset partition to report the results on the Merged/Balance subset which has a balanced proportion of different retrieval strategies and positive/negative samples. For our task, we only use the test set data for the cross-dataset evaluation.
4.1.2. Evaluation Metrics.
We treat the multimodal fake news detection problem as a binary classification task. Following previous works (Shao et al., 2023a; Nan et al., 2022), we apply the Area Under the Receiver Operating Characteristic Curve (AUC), Equal Error Rate (EER), the Accuracy Score (ACC) as our evaluation metrics.
4.1.3. Baselines.
The proposed FKA-Owl is compared with the following strong baseline models: 1) PandaGPT (Su et al., 2023): The off-the-shelf PandaGPT model effectively aligns visual features with the text space of LLMs, enabling it to perform complex multimodal tasks in a zero-shot manner. 2) PandaGPT+SPT: This model integrates PandaGPT with soft prompt tuning (Lester et al., 2021) by using the predefined question prompt along with learnable continuous vectors to fine-tune the LVLM during the instruction tuning phase. 3) HAMMER (Shao et al., 2023a): HAMMER employs two unimodal encoders to encode image and text embeddings with contrastive learning for alignment, and then summarize multimodal information through the multimodal aggregator.
Table 3. Multiple-domain performance (%) comparison of baseline models on DGM4 dataset. Specifically, we use two subsets from the identical country for training and the remaining subsets for testing. SPT denotes the utilization of soft prompt tuning. The better results in each group are in boldface.
@@ 4.1. Experimental Setup
FKA-Owl 75.17 33.35 71.42 75.15 32.85 70.52 81.06 26.21 75.26 85.88 22.41 78.87
4.1.4. Implementation Details.
We utilize the visual encoder and text encoder sourced from ImageBind-Huge (Girdhar et al., 2023) as the backbone for our image and text feature extractors. Moreover, we employ Vicuna-7B (Chiang et al., 2023) as the inferential LLM, connected through a linear layer. The model is initialized from the instruction-tuned checkpoint provided by PandaGPT. All training images are resized to 224 × 224 and subjected to random horizontal flipping, as well as random perturbation techniques such as JPEG compression and Gaussian Blur following (Shao et al., 2023a). The multi-level patch features are extracted from the 8th, 16th, 24th, and 32nd layers of the image encoder. We set the learning rate as 1.5e-5 with a batch size of 16 and a maximum of 10 epochs when trained on the BBC subset. Linear warm-up and the one-cycle cosine learning schedule are adopted. All experiments are conducted on four NVIDIA GeForce 3090 GPUs with PyTorch. More details about hyperparameter settings are provided in the Supplementary Materials.
Table 4. Cross-dataset performance (%) comparison of baseline models on the NewsCLIPpings dataset when trained on DGM4 dataset. Specifically, we use multiple subsets from the DGM4 dataset for training and the NewsCLIPpings dataset for testing. The better results in each group are in boldface.
Test Method BBC & Guardian USA & Wash. Average
@@
NewsCLIPpings PandaGPT+SPT 50.84 49.50 50.22 50.62 49.55 50.33 50.73 49.53 50.28
HAMMER 55.15 45.95 53.92 57.00 44.85 54.56 56.07 45.40 54.24
FKA-Owl 56.41 45.36 54.46 59.03 43.31 56.22 57.72 44.34 55.34
Table 5. Ablation study of component modules. We evaluate the AUC (in %), EER (in %), and ACC (in %) of variant models on the remaining three subsets when trained on the BBC subset. ML: the extraction of multi-level patch features in the cross-modal reasoning (CR) module. DB: the extraction of dual-branch artifact features in the visual-artifact (VA) Localization module. Avg. denotes the mean value on the three testing subsets.
@@
ML&CM Reasoning DB&VA Localization Guardian USA Wash. Avg.
@@
4.2. Performance Comparison
We evaluate the cross-domain performance of our FKA-Owl with baselines in single-domain, multiple-domain, and cross-dataset settings respectively.
4.2.1. Single-domain Setting.
Table 2 presents the performance of our method and other baseline models in the challenging scenario where a single domain is available. We randomly select one subset of the DGM4 dataset as the source domain for training and the remaining subsets as the target domains for testing. From the results, we make the following observations:
  • The performance of the existing MFND method drops significantly when tested on the unknown subsets, which verifies the existence of domain shift caused by the deviation in propagation contents.
  • FKA-Owl yields substantial improvement on recent LVLMs, PandaGPT, and PandaGPT with soft prompt tuning, in both intra-domain and cross-domain testing. Such huge improvement demonstrates the effectiveness of forgery-specific knowledge augmentation in our framework.
  • Compared with the state-of-the-art method, HAMMER, our approach shows more remarkable improvement in cross-domain testing. For instance, for models trained on the BBC subset, FKA-Owl achieves a 7.7% increase in AUC when testing on the Washington Post subset. This may be credited to the effective utilization of inherent world knowledge from LVLMs in mitigating distribution discrepancies. The combination of forgery-specific knowledge and world knowledge facilitates profound manipulation reasoning in FKA-Owl.
4.2.2. Multiple-domain Setting.
The inclusion of domestic news such as the BBC and The Guardian from British, as well as USA TODAY and The Washington Post from America, increases dataset diversity in practical scenarios. We select two subsets from identical countries for training and the remaining two for testing. The results are summarized in Table 3. Our FKA-Owl exhibits significant superiority over both PandaGPT using soft prompt tuning and HAMMER by a large margin. This reveals the effectiveness of our framework in instance-wise domain generalization guided by world knowledge derived from LVLMs, even when jointly learning multiple source domains.
4.2.3. Cross-dataset Setting.
To represent the contextual diversity inherent in real-world multimodal fake news, we select multiple subsets from DGM4 belonging to the same country for training and report the performance on the NewsCLIPpings. As shown in Table 4, the performance of all methods notably decreases when performing cross-dataset testing, which implies that the distribution of different datasets does differ. Furthermore, we can observe that incorporating forgery-specific knowledge into LVLMs markedly improves their ability to detect fake news and consistently surpasses the SOTA model HAMMER, reflecting the generalizability of our approach.
Table 6. Ablation study of prompt strategies. AUC (in %) of variant models is reported on the remaining three subsets when trained on the BBC subset. SPT denotes soft prompt tuning, whereas CAH refers to candidate answer heuristics.
@@ 4.2. Performance Comparison
FKA-Owl 84.95 73.10 74.81 77.62
4.3. Ablation Study
We perform several ablation experiments to explore the necessity of the proposed component modules, prompt strategies, and LVLMs knowledge respectively, and analysis of potential module choices. In the following experiments, all ablation results are evaluated on the remaining three subsets when trained on the BBC subset.
4.3.1. The Effect of the Component Modules.
In Table 5, we conduct a comprehensive ablation study on the proposed component modules to verify their effectiveness. The first row of Table 5 shows our baseline model that only performs soft prompt tuning, achieving an average AUC of 52.5%. Based on this baseline, we further introduce two separate modules: a multi-level cross-modal (ML&CM) reasoning module and a dual-branch visual-artifact (DB&VA) localization module, with 21.55% and 20.16% improvement in average AUC, respectively. This implies that forgery-specific knowledge augmentation is indispensable for our framework. Comparatively, our model with complete two modules obtains the best performance increasing by 25.12%, indicating the effectiveness and complementarity of these two modules. Moreover, we test the performance of FKA-Owl removing the multi-level features (w/o ML), and FKA-Owl removing any one of the dual-branch features (w/o DB). These variant models lack rich features to represent forgery-specific knowledge leading to a great decrease in cross-domain performance.
4.3.2. The Effect of the Prompt Strategies.